OpenStudy (anonymous):

the table summarizes the scoring of a football game between Team A and Team B. A touchdown (TD) is worth 6 points, a field goal (FG) is worth 3 points, a safety (S) is worth 2 points, and a point after touchdown (PAT) is worth 1 point. Use matrix multiplication to find the final score. TD FG S PAT Team A 3 2 1 0 Team B 4 4 1 2 help is appreciated anyone to help first gets a medal

OpenStudy (anonymous):

I am not sure about your information, is it TD FG S PAT Team A 3 2 1 0 Team B 4 4 1 2

OpenStudy (anonymous):

\[\left[\begin{matrix}3 & 2&1&0 \\ 4 & 4&1&2\end{matrix}\right]\left(\begin{matrix}6 \\ 3\\2\\1\end{matrix}\right)\]

OpenStudy (anonymous):

multiply them you get a 2x1 matrix, top is score of A, bottom is score of B That's it